When a pendulum of length 10 cm has swung so that πœƒ is the radian measure of the angle formed by the pendulum and a vertical line, then if β„Ž(πœƒ) cm is the vertical height of the end of the pendulum above its lowest position, β„Ž(πœƒ) = 20 𝑠𝑖𝑛^2 πœƒ/2 . Find the instantaneous rate of change of β„Ž(πœƒ) with respect to πœƒ when (a) πœƒ = πœ‹/3 and (b) πœƒ = πœ‹/2


1
Expert's answer
2021-12-13T17:02:03-0500

h = L (1-cos T) geometry exact


cos 2A = 1 - 2 sin^2 A identity

so

cos T = 1 - 2 sin^2 (T/2)


h = L [ 1 - (1 -2 sin^2 (T/2) ]


= L [2 sin^2 (T/2) ]


= 20 sin^2 (T/2)